OpenStudy (anonymous):

Find the slope and y-intercept of the graph of the equation. x − 6y = 0

hartnn (hartnn):

can you isolate y from x-6y=0 ?

hartnn (hartnn):

add 6y on both sides, what u get?

OpenStudy (anonymous):

x=6y

hartnn (hartnn):

so, y= ... ?

OpenStudy (anonymous):

0?

hartnn (hartnn):

x=6y so y = (1/6)x comparing this with y=mx+c m=slope = 1/6 c= y-intercept = 0